OpenStudy (anonymous):

How would you find the domain and range of g(x)=x^2-6x+4

OpenStudy (anonymous):

domain is the set of values of x for which this function has sence. range is a set of values of g(x) for values of x of the domain

OpenStudy (anonymous):

so how would u find out tho like what formula would you use

OpenStudy (anonymous):

there is no formula for this

OpenStudy (anonymous):

this function has sence for all values of x, so domain is all R

OpenStudy (anonymous):

you really look at the formula given and see if you can make the denominator zero for instance. If you can then this point does not make sense.

OpenStudy (anonymous):

that doesn't apply in this case but is an example of where a domain is incorrect

OpenStudy (anonymous):

i still very confused

OpenStudy (anonymous):

look at the tan function does it have a domain for all values of theta?

OpenStudy (anonymous):

so, domain is all R. To find the range, notice that this is a parabola opened up. So find the vertex to get the minimum value. Range will be (min value, infinity)

OpenStudy (anonymous):

you can see tan x goes vertical every now and then so the domain of tanx is undefined at pi/2 for instance do you see

OpenStudy (anonymous):

it has no value over all real numbers

OpenStudy (anonymous):

to find vertex use this formula for x-coordinate: -b/2a which in this cace gives x=3 so g(3)=-5 so domain: R range: (-5,infinity)

OpenStudy (anonymous):

Another example y=4/x is undefined at x=0 because of division by zero so x=0 is not in the domain of this function?

OpenStudy (anonymous):

Algebraically speaking, how would one find the range and domain of said problem g(x)=x^2-6x+4

OpenStudy (anonymous):

i just explained you that

OpenStudy (anonymous):

 this is, more or less the graph of this function